# Together We Go Further: LLMs and IDE Static Analysis for Extract Method Refactoring

### v1

@@ 4. Evaluation
    How effective are LLMs at generating refactoring suggestions? Answering this is important because this is the first paper that uses LLMs for a novel usage scenario. We conducted a quantitative analysis to measure the effectiveness of LLMs in generating Extract Method refactoring suggestions using three of the latest and most extensive LLMs available.
@@
    How do refactoring suggestions change with different LLM parameters? This is important because we are discovering best practices that others can use for unleashing the full potential of LLMs for refactoring tasks. Moreover, studying how an LLM adjusts refactoring suggestions to different LLM parameters is crucial for tool integration. To answer this question, we conduct a sensitivity analysis to assess the quality of refactoring suggestions in terms of the recall rate among top-5 suggestions (Recall@5) across various LLM parameters and the contributions brought by our enhancements and rankings.
@@
    How effective is EM-Assist in providing refactoring recommendations over existing approaches? To quantitatively evaluate this, we conducted a baseline comparison with 6 other tools that employ static analysis or machine learning techniques to suggest Extract Method refactorings. We compare these tools in terms of Recall@5 using the same corpus that they previously used.
@@
    How useful are the provided recommendations to developers? As we want to advance the science and practice of refactoring, and because EM-Assist is a code renovation tool that developers use interactively, it is important to evaluate whether EM-Assist makes suggestions that developers accept. We employ firehouse surveys with professional developers from our collaborating enterprises, focusing on newly created long methods they committed into code repositories. Then, we report their responses regarding the refactoring recommendations proposed by EM-Assist.
4.1. Datasets
To answer our research questions, we use two kinds of datasets: the “Community Corpora” previously used by other researchers working in this field, and the “Extended Corpus” which we released to address the limitations of previous corpora and to increase its size. Within the Community corpora, we distinguish between two groups: Community Corpus-A and Community Corpus-B, mutually exclusive, each built and maintained by distinct groups of researchers.
@@ 4.1. Datasets
    Community Corpus-A: consists of 122 Java methods and their corresponding Extract Method refactorings collected from five open-source repositories: MyWebMart, SelfPlanner, WikiDev, JHotDraw, and JUnit. This dataset previously served as the foundation for evaluating various state-of-the-art Extract Method refactoring automation tools, including JExtract (Silva et al., 2015), JDeodorant (Tsantalis and Chatzigeorgiou, 2011), SEMI (Charalampidou et al., 2017), GEMS (Xu et al., 2017), and REMS (Cui et al., 2023). While Community Corpus-A is the most widely used by other researchers, it might be interpreted as being subjective. It contains a mix of (i) hypothetical refactoring scenarios in which human experts determined which statements should be extracted from host methods in open source projects and (ii) synthetic refactorings that the corpus designers created by inlining existing methods and then using these expanded host methods to extract the same methods that were previously inlined. While these refactorings are realistic, they are not actual refactorings that open-source developers performed in their code. Thus, we use two other corpora that address this limitation.
@@
    Community Corpus-B: Silva et al. Silva et al. (2016) maintain an active corpus containing 448 Java methods, each accompanied by its respective extract method refactorings that open-source developers actually performed. Importantly, this corpus spans all three categories of refactorings, as classified by Cossette and Walker Cossette and Walker (2012): automatable, non-automatable, and partially automatable by tools that try to replicate a real-world code change scenario. Non-automatable changes are those in which the developers combined refactorings with significant additional feature implementations (either in the host method or in the extracted method) in the same commit. Partially automatable changes are those in which developers made minor changes, such as renaming variables. Automatable changes represent “clean” commits in which the developers made a single type of change, i.e., performed Extract Method refactoring. We chose the latter category as our focus is on replicating the exact development scenarios where our tool only performs refactorings, it does not expand the code functionality. This resulted in 154 replicable refactorings.
  3. (iii)
    Extended Corpus: To enhance the robustness of our evaluation with a sizable oracle of actual refactorings performed by developers, we constructed Extended Corpus. To create it, we employed RefactoringMiner (Tsantalis et al., 2022), the state-of-the-art tool for mining refactorings from commits, with a reported precision of 99.8% and recall of 95.8% for detecting Extract Method. We ran it on highly regarded open-source repositories: IntelliJ Community Edition (JetBrains, 2023b) and CoreNLP (JetBrains, 2023a). After filtering to remove refactoring commits that mixed feature additions (the one-liners and the extracted methods whose body overlapped a large proportion of the host method), we retain 2,849 Extract Methods from these repositories.
4.2. RQ1: Effectiveness of LLMs
@@
Numerous LLMs have been developed by both open-source communities and proprietary companies (Scao et al., 2022; GoogleAI, 2023; OpenAI, 2023; Anthropic, 2023; Meta, 2023; Falcon, 2023). Among them, (i) PaLM (GoogleAI, 2023), (ii) GPT-3.5 (Brown et al., 2020), and (iii) GPT-4 (OpenAI, 2023) are well-known and the largest LLMs, developed by Google and OpenAI, and we use them for the experiments. While these models were not developed specifically for code refactoring tasks but are versatile, we want to determine their effectiveness in generating Extract Method refactoring suggestions.
@@
Figure 3. The capabilities of LLMs in generating refactoring suggestions. The plots show the number of suggestions per host method (notice the exponential scale)
4.2.1. Subject Systems and Experimental Setup
@@
Step 1: We used Community Corpus-A and then prompted the LLM with each method to generate refactoring suggestions. To maximize the capabilities of LLMs to generate suggestions, we employed an iterative approach by adjusting the temperature parameter during LLM prompting, as explained in Section 3.1.2. Specifically, we conducted fixed-point iterations, repeatedly prompting the LLM until it no longer generated new suggestions for each temperature value from the set {0, 0.2, 0.4, 0.6, 0.8, 1.0, 1.2}. This iterative strategy allowed the LLM to create refactorings with varying levels of randomness and expand its search with each subsequent response to the prompt. Then, we analyze the quality of the suggestions by quantitatively studying invalid suggestions (Definition 3.4), not useful suggestions (Definition 3.5), and applicable suggestions (Definition 3.6).
@@ 4.2.2. Results
In Figure 3, box and violin plots show the distribution of suggestions per host method. Starting from the left, total suggestions, then those that are invalid, not useful, and finally, the applicable suggestions generated by each LLM. The data reveals that all three studied LLMs excel in generating refactoring suggestions, with GPT-3.5, GPT-4, and PaLM averaging 20, 11, and 5 refactoring suggestions per host method, respectively. However, they also produced invalid suggestions with averages of 7, 5, and 3, respectively, and not useful suggestions with averages of 4, 1, and 1, respectively. Notably, GPT-3.5 yielded the highest number of applicable suggestions, with an average of 8 per host method, compared to 6 and 2 averages for GPT-4 and PaLM, respectively. These results highlight the effectiveness of LLMs in generating refactoring suggestions while also highlighting the need for additional techniques to filter hallucinations and retain only applicable suggestions.
@@
To further bolster the validity of this claim, we selected GPT-3.5, the top-performing LLM in terms of applicable suggestions, and applied it to the Extended Corpus. Our observations, consistent with the findings in Step 1, revealed that GPT-3.5 generated 12,387 refactoring suggestions, of which 7,715 were hallucinations, leaving only 4,672 applicable suggestions. This further underscores the importance of employing post-processing techniques on LLM output, they cannot be used as-is for refactoring tasks.
Both GPT models are prolific at generating Extract Method suggestions, but more than half of the suggestions cannot be applied.
4.3. RQ2 : Sensitivity Analysis of EM-Assist
@@
We determine the optimal values for LLM hyper-parameters so that we can harness the full potential of LLMs when generating Extract Method refactorings. Then we study the contribution of our design decisions on enhancing the refactoring suggestions made by EM-Assist. This is crucial for revealing best practices for using LLMs for refactoring tasks, and it also enables us to quantitatively assess EM-Assist’s advancements over the raw performance of LLMs.
4.3.1. Subject Systems and Experimental Setup
We use two oracle datasets, namely Community Corpus-A and Community Corpus-B. We focus on two key hyper-parameters: the number of iterations ($I$), with values in the set $\{1,2,3,\ldots,10\}$, and Temperature ($T$), for which we use the following values: $\{0,0.2,0.4,0.6,0.8,1.0,1.2\}$. For each combination $(t,i)$, where $t\in T$ and $i\in I$, we compute EM-Assist’s Recall@5 to study how the tool’s ability to suggest refactoring acceptable by humans varies with these settings. We use Recall@5 as the target metric because while it is possible to collect the examples of real-world refactorings from the open-source projects, we cannot judge for sure if other tools’ suggestions are incorrect. Thus, we aim to maximize the tool’s recall while keeping the number of suggestions reasonably low.
We first define an experiment data point: for a given combination of LLM’s Temperature/Iterations, we are replicating one of the refactoring scenarios from the dataset, i.e., we are using EM-Assist to generate applicable suggestions for the given host method from the dataset. Following the practices used by other researchers (Silva et al., 2015; Xu et al., 2017; Cui et al., 2023; Tsantalis and Chatzigeorgiou, 2011), for each experiment data point we evaluate the results using the top-5 applicable suggestions that EM-Assist produced. We compare these to the actual extracted method according to the oracle.
We count a match (or a hit) when a code fragment from any of the top-5 applicable suggestions matches the code fragment from the actual extracted method in the dataset, with an ntolerance. For example, consider an extract method suggestion $e_{i}(n_{i},(s_{i},f_{i}))$ where $s_{i}$ and $f_{i}$ are the start and end line numbers of the suggested code fragment to be extracted, while the oracle specifies $e_{j}(n_{j},(s_{j},f_{j}))$ as the actual refactoring. The ntolerance verifies whether the start and end line numbers of the suggestion are within the start and end lines of the ground truth extracted method, with a tolerance threshold at most nof the length of the ground truth method. Formally, $round(s_{j}(1-len\cdot n/100))\leq s_{i}\leq round(s_{j}(1+len\cdot n/100))$ and $round(f_{j}(1-len\cdot n/100))\leq f_{i}\leq round(f_{j}(1+len\cdot n/100))$ where $len=f_{j}-s_{j}+1$.
Then, Recall@5 is the percentage of hits in top-5 at ntolerance relative to the total number of actual extract methods in the oracle dataset. In other words, we compute the ratio of extract methods that we suggested and are perceived correctly by the oracle, while showing up to five suggestions per example.
We calculate Recall@5 using suggestions at 3% tolerance for each combination $(t,i)$ on a version of EM-Assist with enhancements (see Section 3.4) turned on, and selecting the top-5 suggestions using the heat-popularity ranking $\mathcal{T}_{3}(e_{i})$ (see Section 3.5). Additionally, we compute the metric for EM-Assist, with the enhancements turned off, and using the same heat-popularity ranking function $\mathcal{T}_{3}(e_{i})$. This analysis not only aids in understanding the optimal settings for LLM but also sheds light on the impact of EM-Assist’s design decisions.
@@ 4.3.1. Subject Systems and Experimental Setup
Figure 4. Change of Recall@5 along with Temperature and iterations for the Community Corpus-A
Figure 4. Change of Recall@5 along with Temperature and iterations for the Community Corpus-A
Figure 5. Change of Recall@5 along with Temperature and iterations for the Community Corpus-B
4.3.2. Results
Figure 5 and Figure 5 show via heatmaps how EM-Assist’s effectiveness varies with LLM Temperature/Iterations for the two corpora. The results show that higher temperature values and more iterations consistently yield superior results. For Community Corpus-A, EM-Assist achieved its highest Recall@5 values at 66.4%, with a temperature setting of 1.2 and 10 iterations. For Community Corpus-B, the best metric value was 35.1%, achieved with Temperature 1.0 and 9 iterations.
We also conducted a sensitivity analysis on the impact of EM-Assist’s design decision on improving its effectiveness. Table 1 shows an ablation study to find out the contributions of various modules from the EM-Assist’s pipeline (see Figure 2). For temperature 1.2 and 10 iterations, we compute Recall@5 after various stages in the EM-Assist’s pipeline, starting from EM-Assist’s raw performance with enhancements (i.e., heuristics) and rankings disabled, all the way to activating all modules. The results show that EM-Assist’s design decisions contributed a maximum improvement of 27 percentage points for 1% tolerance. These results highlight the significant impact of EM-Assist’s design decisions.
With a higher temperature and number of iterations, LLMs produce better suggestions for Extract Method recommendation. Our ranking and heuristic techniques further rise the better suggestions to the top.
Table 1. Ablation Study of EM-Assist showing the contributions of heuristics and ranking
@@ 4.3.2. Results
Method Tolerance 1% Tolerance 2% Tolerance 3%
Random rank, no heuristics 33.6% 35.2% 40.1%
Random rank, with heuristics 37.7% 39.3% 44.2%
EM-Assist- ($\mathcal{T}_{1}(e_{i})$) 59.0% 59.8% 65.5%
EM-Assist- ($\mathcal{T}_{2}(e_{i})$) 60.6% 61.4% 66.4%
EM-Assist- ($\mathcal{T}_{3}(e_{i})$) 60.6% 61.4% 66.4%
4.4. RQ3 : Effectiveness of EM-Assist
@@
Table 2. Evaluation results of EM-Assist with respect to six other tools on Community Corpus-A

### v2

@@ IV Evaluation
    How effective are LLMs at generating refactoring suggestions? Understanding the capability of LLMs to produce EM suggestions is essential for our tool’s effectiveness. Therefore, we conducted a quantitative analysis involving three of the latest LLMs available.
@@
    How do refactoring suggestions change with different LLM parameters? This is important for integrating LLMs into tools and for researchers using LLMs for refactoring. We conduct a sensitivity analysis to assess the quality of these suggestions, focusing on the recall rate for the top-5 suggestions (Recall@5) across various LLM parameters, and examine the effects of our enhancements and ranking methods.
@@
    How effective is EM-Assist in providing refactoring recommendations over existing approaches? To quantitatively evaluate this, we conducted a baseline comparison with 6 other tools that employ static analysis or machine learning techniques to suggest EM refactorings.
@@
    How useful are the provided recommendations to developers? EM-Assist is a code renovation tool that developers use interactively. Thus, it is important to evaluate whether EM-Assist makes suggestions that developers accept. We employ firehouse surveys with professional developers from our collaborating enterprises, focusing on newly created long methods they committed into code repositories.
IV-A Datasets
To answer our research questions, we use two datasets: the “Community Corpora” previously used by other researchers working in this field, and the “Extended Corpus” which we released to address the limitations of previous corpora and to increase its size.
@@ IV-A Datasets
    Community Corpus: consists of 122 Java methods and their corresponding EM refactorings using five open-source repositories: MyWebMart, SelfPlanner, WikiDev, JHotDraw, and JUnit. This dataset previously served as the foundation for evaluating various state-of-the-art EM refactoring automation tools, including JExtract [27], JDeodorant [1], SEMI [8], GEMS [11], and REMS [12]. The Community Corpus, although extensively used by researchers, can be seen as subjective. It combines (i) theoretical refactoring scenarios assessed by third-party (usually students) who are not familiar with the open-source projects, and (ii) synthetic refactorings devised by corpus designers, based on inlining existing methods followed by extracting these same methods from the expanded host methods. While these refactorings are realistic, they are not actual refactorings that developers performed. Thus, we created another corpora that addresses these limitations.
@@
    Extended Corpus: To further strengthen our evaluation’s robustness with a broad oracle of actual refactorings by developers and ensure more generalizable results, we constructed Extended Corpus. To create it, we employed RefactoringMiner [6], the state-of-the-art tool for mining refactorings from commits, with a reported precision of 99.8% and recall of 95.8% for detecting EM. We ran RefactoringMiner on 12 open-source repositories, including notable projects like CoreNLP, Guava, and Gson, which span a wide range of domains such as machine learning, database management, data handling, and web technologies. We filtered out refactoring commits that include one-liners and extracted methods whose bodies overlapped significantly with the host method. Additionally, we filtered out non-automatable [43] refactorings, such as those entailing feature additions. For this purpose, we leveraged the state of the practice IDE for EM, IntelliJ IDEA, using start and end line numbers of the extracted code fragment provided by RefactoringMiner to determine whether IntelliJ IDEA could automatically extract the code fragment identified.
    Consequently, we retained 1752 EMs from these repositories. We demonstrate that this corpus is representative in terms of the length of the host method, cyclomatic complexity [44] of the host method, and Halstead metric difficulty [45] of the host method, as shown in Figure 3.
    Figure 3: Metrics for host methods in our corpus
    IV-B Evaluation Metric: Recall@n with mtolerance
    Previous researchers [27, 11, 12, 1] have designed metrics to evaluate EM suggestion tools. In their evaluations, they analyze the top-n ranked suggestions generated by their tools at various $m\%$ tolerance levels. We also employ Recall@$n$ with $m\%$ tolerance to evaluate EM-Assist. First, to explain $m\%$ tolerance level, let us examine a specific instance. Consider an extract method suggestion $e_{i}(n_{i},(s_{i},e_{i}))$ where $s_{i}$ and $e_{i}$ are the start and end line numbers of the suggested code fragment to be extracted, while the oracle specifies $e_{j}(n_{j},(s_{j},e_{j}))$ as the actual refactoring. The $m\%$ tolerance verifies whether the start and end line numbers of the suggestion are within the start and end lines of the ground truth extracted method, with a tolerance threshold at most $m\%$ of the length of the ground truth method, $L_{i}$. Formally, $|f_{i}-f_{j}|+|s_{i}-s_{j}|\leq n/100*L_{i}$. Next, Recall@n is calculated as the percentage of host methods for which at least one of the top-n EM suggestions matches the refactoring specified in the oracle, given $m\%$ tolerance level.
    Given the non-deterministic nature of LLM outputs, recall rates for EM-Assist may vary across invocations. However, the impact of this on EM-Assist is minimal since we invoke the LLM multiple times. Despite the minimal effect, during the experiments, when we report the recall rate for EM-Assist, we repeat the experiments $\mathcal{N}$ times and consider a distribution of recall rates, then report the mean and standard deviation of these values. While a larger $\mathcal{N}$ is preferable, we use $\mathcal{N}$ = 30, following the best practices suggested by Andrea et al.[46].
IV-C RQ1: Effectiveness of LLMs
@@
Numerous LLMs have been developed by both open-source communities and proprietary companies [47, 22, 21, 48, 49, 50]. Among them, (i) PaLM [22], (ii) GPT-3.5 [32], and (iii) GPT-4 [21] are well-known and the largest LLMs, developed by Google and OpenAI, and we use them for the experiments. While these models were not developed specifically for refactoring, they are versatile. We determine their effectiveness in generating EM refactoring suggestions.
@@
Figure 4: The capabilities of LLMs in generating refactoring suggestions. The plots show the number of suggestions per host method (notice the exponential scale)
IV-C1 Subject Systems and Experimental Setup
@@
Step 1: We used Community Corpus and then prompted the LLM with each method to generate refactoring suggestions. To maximize the capabilities of LLMs to generate suggestions, we employed an iterative approach by adjusting the temperature parameter during LLM prompting, as explained in Section III-A2. Specifically, we conducted fixed-point iterations, repeatedly prompting the LLM until it no longer generated new suggestions for each temperature value from the set {0, 0.2, 0.4, 0.6, 0.8, 1.0, 1.2}. This iterative strategy allowed the LLM to create refactorings with varying levels of randomness and expand its search with each subsequent response to the prompt. Then, we analyze the quality of the suggestions by quantitatively studying invalid suggestions (Definition III.4), and not useful suggestions (Definition III.5).
@@ IV-C2 Results
In Figure 4, box and violin plots show the distribution of suggestions per host method. Starting from the left, total suggestions, then those that are invalid, not useful, and finally, the useful suggestions generated by each LLM. The data reveals that all three studied LLMs excel in generating refactoring suggestions, with GPT-3.5, GPT-4, and PaLM averaging 27, 14, and 7 refactoring suggestions per host method, respectively. However, they also produced invalid suggestions with averages of 12, 6, and 3, respectively, and not useful suggestions with averages of 4, 1, and 0, respectively. Notably, GPT-3.5 yielded the highest number of useful suggestions, with an average of 10 per host method, compared to 7 and 2 averages for GPT-4 and PaLM, respectively. These results highlight the effectiveness of LLMs in generating refactoring suggestions while also highlighting the need for additional techniques to select only useful suggestions.
@@
To further bolster the validity of this claim, we selected GPT-3.5, the top-performing LLM in terms of applicable suggestions, and applied it to the Extended Corpus. Our observations, consistent with the findings in Step 1, revealed that GPT-3.5 generated total 22741 refactoring suggestions, of which 17356 were hallucinations (i.e., invalid and non-useful), leaving only 5385 (23.7%) useful suggestions. This underscores the importance of employing post-processing techniques on LLMs’ output, they cannot be used as-is for refactoring tasks.
LLMs excel at generating EM suggestions, yet only 23.7% of these suggestions are useful.
IV-D RQ2 : Sensitivity Analysis of EM-Assist
@@
We determine the optimal values for LLM hyper-parameters so that we can harness the full potential of LLMs when generating EM refactorings. Then we study the contribution of our design decisions on enhancing the refactoring suggestions made by EM-Assist. This is crucial for revealing best practices for using LLMs for refactoring tasks, and it also enables us to quantitatively assess EM-Assist’s advancements over the raw performance of LLMs.
IV-D1 Subject Systems and Experimental Setup
We use our oracle dataset, Community Corpus, to analyze the output of EM-Assist by examining the effects of specific Temperature values ($t\in{0,0.2,0.4,0.6,0.8,1.0,1.2}$) and iteration numbers ($i\in{1,2,\ldots,10}$). For each parameter combination $(t,i)$, we invoke EM-Assist to generate extract method suggestions for a host method within the dataset. In line with the methodologies adopted by previous researchers [27, 11, 12, 1] for assessing earlier EM automation tools, we evaluate each parameter setting by calculating Recall@5 with a 3% tolerance level. Here, to identify the optimal parameter values ($t$, $i$), we analyze a distribution of recall values (as explained in Section IV-B) and select the mean recall. Then, we assess the recall both with and without EM-Assist’s enhancements to the LLM suggestions.
@@ IV-D1 Subject Systems and Experimental Setup
Figure 5: Change of Recall@5 along with Temperature and iterations for the Community Corpus
TABLE I: Ablation Study: Mean and Standard Deviation of Recall Values for different stages in EM-Assist’s pipeline.
@@
Method Tolerance Tolerance Tolerance
1% 2% 3%
Random rank, no heuristics $32.9\%\pm 2.3$ $34.2\%\pm 2.3$ $37.4\%\pm 2.4$
Random rank, with heuristics $44\%\pm 2.8$ $44.8\%\pm 2.9$ $49.6\%\pm 2.9$
Ranking and heuristics 57.6% $\pm$ 1.6 58.4% $\pm$ 1.6 63% $\pm$ 1.7
IV-D2 Results
Figure 5 shows via heatmaps how EM-Assist’s mean recall varies with LLM Temperature/Iterations. The results show that higher temperature values and more iterations consistently yield superior results. Even though LLM randomness increases with higher temperature values, potentially leading to a greater variety of suggestions and possibly more hallucinations with more iterations, the enhancements implemented in EM-Assist effectively refine and rank these suggestions in such a manner that its recall improves alongside these parameters. Notably, EM-Assist recorded its peak Recall@5 scores at 63%, with the configuration set to a temperature of 1.2 and 10 iterations, establishing these parameters as the optimal settings for EM-Assist’s applications.
To further study the contributions of various components within the EM-Assist’s pipeline (see Figure 2), we conducted an ablation study. We structured this into three distinct phases: (i) In the initial phase, we processed the raw output from the LLM without any heuristic enhancements and randomly selected five candidates. (ii) We considered the enhanced output with heuristics (see Section III-D) and arbitrarily choose five suggestions after enhancement. (iii) We took into account the enhanced output and applied the tool’s ranking mechanism. These stages were instrumental in understanding the individual contributions of the various modules to the overall performance. As shown in Table I, the results emphasized that significant improvements were observed across all tolerance levels, with a notable enhancement at the 3% tolerance level. Specifically, the baseline mean recall of the LLM was recorded at 37.4%. Through our design decisions, this figure was elevated to 63%. This enhancement vividly illustrates the substantial influence of the strategic decisions embedded in our tool’s design.
With higher temperature (1.2) and more iterations (10), LLMs produce better suggestions for EMs. Our ranking and heuristics further amplified the quality of the suggestions, achieving an uplift of up to 26 percentage points.
IV-E RQ3 : Effectiveness of EM-Assist
@@
IV-E1 Subject systems and Experimental Setup
We selected six tools that are representative of a wide array of techniques, including four (JDeodorant [1], JExtract [27], SEMI [8], and LiveRef [29]) that use static analysis-based rules and software quality metrics, and two (REMS [12] and GEMS [11]) that use prediction models based on machine learning techniques.
We first employ Community Corpus to evaluate the effectiveness of the selected tools. Following practices employed by the authors of other tools, we evaluate top-5 suggestions generated by the tools, calculating mean Recall@5 at tolerance levels of 1%, 2%, and 3% (for details on evaluation metrics refer to Section IV-B). For EM-Assist, we compute the metric for the best-performing hyper-parameters. For four tools (GEMS, JDeodorant, JExtract, SEMI), we reuse the evaluation results reported by other researchers [11] because they used the same dataset and they defined the same metric formulation as we do. For two tools (REMS and LiveRef), we had to run them ourselves: LiveRef had not previously been executed on Community Corpus, and REMS computes recall differently than any of the other tools. Notice that we contacted the authors and had extensive conversations with them to fully understand how to replicate their results best and how to use their tools so they perform the best.
Second, in order to bolster the robustness of our evaluation, we expand our evaluation to Extended Corpus, which replicates 1752 EM refactorings that took place in open-source projects. In this extended evaluation, we compare directly with JExtract. During our evaluation on Community Corpus, we noticed that GEMS and JExtract had the highest recall rates following EM-Assist, and surpassing all other tools. Their recall rates were remarkably similar, and given the extensive scale of the experiments and the time considerations, we opted for JExtract to execute and benchmark the recall rates against EM-Assist, using a Recall@5 metric at 3% tolerance.
IV-E2 Results
Table II shows the comparative effectiveness of EM-Assist in relation to the other six tools. We tame the non-determinism of LLMs as detailed in Section IV-B. Among the tools evaluated, GEMS ranked second highest, with JExtract closely behind. EM-Assist achieves superior performance with higher Recall@5 scores at 1% and 3% tolerance levels, recording values of 57.6% and 63%, respectively.
To further strengthen the validity of our results, we applied both EM-Assist and JExtract on the Extended Corpus that includes 1752 actual refactorings from open-source projects. For both tools, we calculated the recall for top-5 suggestions with a 3% tolerance value. As shown in Table III, JExtract achieved recall of 39.4%, while EM-Assist’s mean recall was 53.4%, which is 35% improvement over the baseline. This shows that although EM-Assist’s advantage is modest when evaluated on a synthetic Community Corpus, its performance notably surpasses the baseline when it comes to replicating refactorings actually performed by developers.
To address the non-determinism of LLM outputs, we statistically analyzed the recall values of EM-Assist and JExtract over the Extended Corpus. We created a distribution of recall values for EM-Assist and then employed the Wilcoxon Signed-Rank Test to compare this distribution against the recall values of JExtract. The test rejected the null hypothesis, indicating a statistically significant difference in recall. Following this, we used the Hodges-Lehman estimator on each combination of distributions to quantify the difference in recall rates. The resultant value was 12.4 percentage points, suggesting a notable difference in performance between the two tools. This shows that EM-Assist’s improvements over the baseline are statistically significant.
TABLE II: Evaluation results of EM-Assist with respect to six other tools on Community Corpus